OpenStudy (pengy120):

i need help on these three questions (simplifying) 1) 3√20-2√45 2) 5√18+4√32 3) √12-7√75

OpenStudy (tkhunny):

Have you considered hunting for perfect squares? 20 = 4*5 45 = 9*5 18 = 9*2 Like that.

OpenStudy (pengy120):

oooooooh okay, thank you!

OpenStudy (pengy120):

when i split them into perfect squares would i put the three on both sides? like 3√20=3√4 and 3√5?

OpenStudy (tkhunny):

\(3\sqrt{20} = 3\sqrt{4\cdot 5} = 3\sqrt{4}\sqrt{5} = 3\cdot 2\sqrt{5} = 6\sqrt{5}\)

OpenStudy (pengy120):

thanks